OpenStudy (anonymous):

If N = {all real numbers} and subset S = {multiples of 4 less than zero}, what is S′? {multiples of 4 greater than zero} {even numbers less than zero} {odd numbers less than zero} {all real numbers excluding multiples of 4 less than zero}

OpenStudy (zzr0ck3r):

S' = N\S?

OpenStudy (anonymous):

@shahbaaz786 \(\bf S'\) is the complement of \(\bf S \) then \(\bf S'\) contains all elements which S doesn't contain. Hence:\[\bf \left\{ S \subset N \ | N = \mathbb{R}, \ S=4k<0 \ for \ k < 0 \right\} \implies S' = N-S=N / S\] @zzr0ck3r is right. Note that my "/" should really be "\" to denote a relative set but the equation editor doesn't allow me to do that for some reason...

OpenStudy (nory):

Isn't S' the complement of S?

OpenStudy (zzr0ck3r):

it is, and here s is a subset of universe N = R so R\S remove S from R

OpenStudy (nory):

So therefore, it would be all elements of R that are not contained in S.

OpenStudy (anonymous):

yup, that's right.

OpenStudy (nory):

Which is, all real numbers that are not multiples of 4 less than 0. Does that make sense?

OpenStudy (zzr0ck3r):

its important to state the universe of shich S is being complimented with

OpenStudy (zzr0ck3r):

Z\S and R\S are two different things

OpenStudy (nory):

True...

OpenStudy (zzr0ck3r):

IOW S' depends on the universe S is a subset of

OpenStudy (anonymous):

so its D ?

OpenStudy (anonymous):

two very different things..

OpenStudy (nory):

Well, if it were Z then instead of real numbers it would be integers, et cetera. Yes I think it is D.

OpenStudy (anonymous):

@shahbaaz786 correct.

OpenStudy (zzr0ck3r):

correct

OpenStudy (zzr0ck3r):

its trivial, but we need to know that S was in N=R

OpenStudy (anonymous):

@zzr0ck3r we already know that...

OpenStudy (zzr0ck3r):

p.s. your book is lame for calling R = N

OpenStudy (anonymous):

and i agree lol since the N can easily be confused with the natural numbers... @zzr0ck3r